Hi
George set off at noon driving at 80km per hour and reached his destination at 10pm. Calvin set off at noon as well and reached the same destination at 8pm.
How far was George from his destination when he and Calvin were 120km apart.
Thanks

Answer by ikleyn(52925)

.
George set off at noon driving at 80km per hour and reached his destination at 10pm.
Calvin set off at noon as well and reached the same destination at 8pm.
How far was George from his destination when he and Calvin were 120km apart ?
~~~~~~~~~~~~~~~~~

The distance to the destination was 80*10 = 800 kilometers.  (10 hours is the George's travel time).


Hence, the Calvin' speed (rate) was  800/8 = 100 km/h.


Thus, their separation rate was  100 km/h - 80 km/h = 20 km/h.


It means that they were 120 km apart in  120/20 = 6 hours after start (after the noon).


At this time moment, George was  80*(10-6) = 80*4 = 320 kilometers from his destination.    ANSWER

Solved.